OpenStudy (anonymous):

If a stone is tossed from the top of a 270 meter building, the height of the stone as a function of time is given by h(t) = -9.8t2 – 10t + 270, where t is in seconds, and height is in meters. After how many seconds will the stone hit the ground? Round to the nearest hundredth’s place; include units in your answer. (Points : 20)

OpenStudy (saifoo.khan):

270 = -9.8t2 – 10t + 270 Solve.

OpenStudy (saifoo.khan):

-9.8t2 – 10t = 0

OpenStudy (lalaly):

i answerd this question http://openstudy.com/users/lalaly#/users/lalaly/updates/4e6871930b8b4a2b95c92574 jus instead of 310 change it to 270

OpenStudy (anonymous):

\[ h(t) = -9.8t^2 – 10t + 270\]Substitute h with 270\[270 = -9.8t^2 – 10t + 270\]\[ -9.8t^2 – 10t =0\]Solve for t\[t(-9.8t – 10) =0\] \[t=0\]or\[-9.8t-10=0\]\\[t=\frac{10}{-9.8}\] Both solutions are rejected.